OpenStudy (anonymous):

P=KT/v when T=30,P=3 ,k=0.5 how do you solve for V?

OpenStudy (anonymous):

P = KT/v v = KT/P v = (0.5)(30)/(3) = 5

OpenStudy (anonymous):

How u get five

OpenStudy (anonymous):

well 0.5 times 30 is 15 and 15 divided by 3 is 5 that's how i got 5

OpenStudy (anonymous):

Can you answer the other question ppplllease

OpenStudy (anonymous):

The one posted after this one